40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 14 22:56:31 2011, kiwamu, Update, Green Locking, Y end PDH lock : UGF at 17 kHz YendOLTF.png
    Reply  Tue Nov 15 09:51:04 2011, Zach, Update, Green Locking, Y end PDH lock : UGF at 17 kHz 
Message ID: 5890     Entry time: Mon Nov 14 22:56:31 2011     Reply to this: 5893
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y end PDH lock : UGF at 17 kHz 

[Tomotada / Kiwamu]

  The open loop transfer function of the Y end PDH loop was remeasured : the UGF was found to be at 17 kHz.

The phase margin at the UGF was about 27 deg.

YendOLTF.png

 

While the measurement we noticed that the modulation onto the laser PZT was too big

and it was creating a big AM on the reflected light with an amplitude of a few mV.

So we put a 20 dB attenuator to decrease the modulations and the reflected light became much quitter.

Also the servo shape formed by Newfocus LB1005 looks too simple : we should have a more sophisticated servo filter (i.e. PDH box!!).

ELOG V3.1.3-